If there is 5.0 g of CO2 gas in a 10.0 L cylinder at 25˚C, what is the gas pressure within the cylinder?


1
Expert's answer
2021-11-05T01:07:08-0400

PV=nRT

P=?

V= 10.0L

T=273+25=298K

R=8.3145

"P×10=\frac{5.0}{44}×8.3145×298"


"P\u00d710=281.56"


"P=28.156Pa"